# Robust Stochastically-Descending Unrolled Networks

### v1

@@ Abstract
Deep unrolling, or unfolding, is an emerging learning-to-optimize method that unrolls a truncated iterative algorithm in the layers of a trainable neural network. However, the convergence guarantees and generalizability of the unrolled networks are still open theoretical problems. To tackle these problems, we provide deep unrolled architectures with a stochastic descent nature by imposing descending constraints during training. The descending constraints are forced layer by layer to ensure that each unrolled layer takes, on average, a descent step toward the optimum during training. We theoretically prove that the sequence constructed by the outputs of the unrolled layers is then guaranteed to converge for unseen problems, assuming no distribution shift between training and test problems. We also show that standard unrolling is brittle to perturbations, and our imposed constraints provide the unrolled networks with robustness to additive noise and perturbations. We numerically assess unrolled architectures trained under the proposed constraints in two different applications, including the sparse coding using learnable iterative shrinkage and thresholding algorithm (LISTA) and image inpainting using proximal generative flow (GLOW-Prox), and demonstrate the performance and robustness benefits of the proposed method.
@@ I Introduction
The lack of convergence across the unrolled layers engenders two critical issues despite the superior performance achieved at the last layers. First, the lack of convergence guarantees limits our ability to draw conclusions about the generalizability of these models to unseen problems. Contrarily, standard iterative algorithms are endowed with theoretical convergence guarantees, thereby eliminating any doubts about their generalizability. Second, unrolled networks are brittle to perturbations added to the layers’ outputs due to the lack of convergence guarantees. In contrast, standard iterative algorithms with descending guarantees, like GD, are known to be resilient to perturbations as they would still follow a descending direction in the iterations that follow the one where the perturbation occurs (see Figure 2 (left)). This evidently does not extend to their unrolled counterparts since they lack descending guarantees. Revisiting our example, Figure 2 (middle) confirms this fact as the unrolled optimizer fails to converge to the same point as in Figure 1 (middle) after adding perturbations to one of its layers; in fact, the figure shows that the distance to the optimal at the last layer is larger than it is at initialization. This raises concerns about implementing these models in safety-critical applications.
@@
Previous works have investigated the convergence issue of algorithm unrolling. One approach for providing convergence guarantees, known as safeguarding, has been proposed in [27, 28, 29, 30]. The main idea of safeguarding is to check whether the estimate made by a certain layer, during inference, satisfies some form of descent criteria. If the criteria are not satisfied, the estimate is discarded and replaced with an estimate of the classic iterative algorithm. Another line of work proved the existence of a convergent unrolled network for specific iterative algorithms such as the iterative shrinkage and thresholding algorithm (ISTA) [31, 32] and differentiable linearized ADMM [33]. However, a mechanism that guarantees to find these convergent networks was never provided. To resolve this issue, [32, 34] suggest reducing the size of the search space by learning fewer parameters of the standard algorithm, which limits the network’s expressivity. Furthermore, all these theoretical endeavors are tailored for specific optimization algorithms and do not generalize to other algorithms. This leaves many existing and potential unrolled networks without convergence guarantees.
@@
In this paper, we aim to design unrolled algorithms that are stochastically descending as in Figure 1 (right) and robust to additive perturbations as in Figure 2 (right). On top of that, our goal is to provide theoretical convergence guarantees for unrolled networks regardless of the underlying iterative algorithms. We accomplish this by posing the training problem as a constrained learning problem. Under this constrained formulation of the problem, we force each unrolled layer to take a descent direction, on average. This descent direction need not be the gradient direction, as shown in Figure 1 (right), in order to accelerate convergence even for harder landscapes. To complement our endeavor, we theoretically derive convergence guarantees for unrolled networks trained with our descending constraints and prove exponential rates of convergence. In summary, our contributions are as follows:
@@
  • We empirically show the robustness of these unrolled networks to perturbations in light of two applications: sparse coding (Section IV) and image inpainting (Section V).
@@ II Problem Formulation
The inner problem is an instantiation of finding a minimum of $f(\cdot;{\bf x})$ for a given ${\bf x}$, which we refer to as the optimizee. It, however, needs to be evaluated for all ${\bf x}\in{\cal X}$, which is computationally expensive. The outer problem aims to learn the optimizer $\bm{\Phi}:{\cal X}\rightarrow{\cal Y}$ by minimizing a suitably-chosen loss function $\ell:{\cal Y}\times{\cal Y}\rightarrow\mathbb{R}$. For instance, the toy example in Figure 1 considers a function class $\mathfrak{F}$ that contains functions of the form $f({\bf y};{\bf A},{\bf z})=\|{\bf Ay}-{\bf z}\|_{2}$, which is parameterized by ${\bf A}\in{\cal S}$ and ${\bf z}\in{\cal Z}$ with ${\cal X}\equiv{\cal S}\times{\cal Z}$. The optimizee problem is then to find the stationary point of $f$ for each combination of $\bf A$ and $\bf z$ in $\cal X$. The optimizer $\bm{\Phi}$, in this case, is designed by unrolling GD in $L$ layers, each of which is a multilayer perceptron (MLP) with residual connections, i.e., the $l$th layer’s output is ${\bf y}_{l}={\bf y}_{l-1}-\text{MLP}({\bf y}_{l-1},{\bf A},{\bf z};{\bf W}_{l})$. These MLPs are trained by minimzing the mean-square error (MSE) between the final prediction of $\bm{\Phi}$ and the optimal ${\bf y}^{*}$.
Solving the above bi-level optimization problem is not usually straightforward. However, it reduces to an unconstrained problem if the inner problem either has a closed-form solution or can be solved using an iterative algorithm. For our problems of interest, we assume that there exists an iterative algorithm ${\cal A}$ that finds a (local) minimum ${\bf y}^{*}$ of the optimizee problem. Given algorithm $\cal A$ and, therefore, the optimal solution to the optimizee, the optimizer problem can be then simplified to
@@ III-A Descending Constraints
The initial value ${\bf y}_{0}$ is drawn from a Gaussian distribution ${\cal N}(\mu_{0},\sigma_{0}^{2}{\bf I})$. We also add noise vector ${\bf n}_{l}$, sampled from a Gaussian distribution ${\cal N}(0,\sigma^{2}_{l}{\bf I})$, to the output ${\bf y}_{l}$ before it is fed to the next layer to make the constraints statistically independent. The noise variance $\sigma_{l}^{2}$ decreases over the layers to allow the optimizer to converge.
@@
It is worth noting that the choice between (CI) or (CII) is arbitrary. The former requires the objective function $f$ to be differentiable while the latter depends on the knowledge of the optimal ${\bf y}^{*}$. Yet (CI) can be extended to non-differentiable functions by replacing the gradient with a subgradient of $f$. However, (CI) could be more challenging to satisfy if the optimizee problem is either constrained or non-convex. This is true since it could push the estimates to a non-feasible solution in the first instance or another local minimum than ${\bf y}^{*}$ in the second one.
@@ III-B Probably and Approximately Correct Solution
To solve (CO), we construct the dual problem to leverage CLT [40, Theorem 1]. The dual problem of (CO) is equivalent to finding the saddle point of the Lagrangian function
@@
Solving (DO) and finding the saddle point of $\widehat{\cal L}$ is doable by alternating between minimizing with respect to ${\bf W}$ for a fixed $\bm{\lambda}$ and then maximizing the latter, as described in Algorithm 1. The convergence of this algorithm has been proved in [40, Theorem 2].
@@ III-C Convergence Guarantees
Given are a near-optimal solution of (CO) under constraints (CI), which satisfies (8) with a probability $1-\delta$ and generates a sequence of random variables $\{{\bf y}_{l}\}_{l=1}^{L}$. Then, under Assumption 1, it holds that
@@
$\begin{split}\mathbb{E}\big{[}{\|{\nabla}f({\bf W}_{L})\|}\big{]}\leq(1-\delta)^{L}(&1-{\epsilon})^{L}\ \mathbb{E}{\|{\nabla}f({\bf W}_{0})\|}\\ &+\frac{1}{\epsilon}\left(\zeta(Q,\delta)+\frac{\delta M}{1-\delta}\right).\end{split}$ (10)
@@ IV-A Problem Formulation
$\begin{split}\min_{{\bf y}}\quad&\frac{1}{2}\|{\bf x}-{\bf Dy}\|_{2}^{2}+\alpha\|{\bf y}\|_{1},\end{split}$ (11)
@@ IV-D Numerical Results and Discussions
Performance. Figure 3 compares the performance of LISTA and constrained LISTA on a test dataset. The two models achieve the same performance at their outputs, illustrated by the zero distance achieved at the last layer in both cases. However, their behaviors over the intermediate layers differ significantly from each other. For standard LISTA, the distance to the optimal is almost the same across all these layers, which suggests that there was no significant improvement in the estimates. In contrast, constrained LISTA demonstrates a gradual decrease in the distance to the optimum because of the imposed constraints.
Figure 3: Distance to the optimal solution ${\bf y}^{*}$ of the sparse coding problem (11) across the unrolled layers of LISTA (red) vs constrained LISTA (blue). The distance does not change significantly at the intermediate layers of LISTA, while constrained LISTA makes gradual progressions toward the optimum.
@@
Figure 5: Robustness against additive noise with a standard deviation proportional to the gradient by a factor $\hat{\sigma}$. (Left) Distance to the optimal ${\bf y}^{*}$, (middle) $\ell_{1}$-norm of the unrolled network’s output, and (right) Frequency of satisfying the constraints across the test dataset for different perturbation sizes.
@@
Robustness. To examine the impact of the descending constraints, we show in Figure 5 the performance of LISTA and constrained LISTA after introducing additive perturbations at each unrolled layer. The perturbation is modeled by a noise vector, where each element is normally distributed with zero mean and a standard deviation of $\sigma_{l}=\hat{\sigma}g_{l}$. The figure confirms the robustness of constrained LISTA since the deterioration in both the distance to the optimal and the $\ell_{1}$-norm of the estimate is very slow compared to LISTA. Figure 5 (right) also shows that all the constraints were upheld in the case of constrained LISTA for different values of $\hat{\sigma}$, which confirms the robustness of the model.

### v2

@@ Abstract
Deep unrolling, or unfolding, is an emerging learning-to-optimize method that unrolls a truncated iterative algorithm in the layers of a trainable neural network. However, the convergence guarantees and generalizability of the unrolled networks are still open theoretical problems. To tackle these problems, we provide deep unrolled architectures with a stochastic descent nature by imposing descending constraints during training. The descending constraints are forced layer by layer to ensure that each unrolled layer takes, on average, a descent step toward the optimum during training. We theoretically prove that the sequence constructed by the outputs of the unrolled layers is then guaranteed to converge for in-distribution problems. We then analyze the generalizability to certain out-of-distribution (OOD) shifts in the optimization problems being solved. Our analysis shows that the descending nature imposed by the proposed constraints is transferable under these distribution shifts, subject to a generalization error, thereby providing the unrolled networks with OOD robustness. We numerically assess unrolled architectures trained with the proposed constraints in two different applications, including the sparse coding using learnable iterative shrinkage and thresholding algorithm (LISTA) and image inpainting using proximal generative flow (GLOW-Prox), and demonstrate the performance and robustness advantages of the proposed method.
@@ I Introduction
The lack of convergence across the unrolled layers engenders a critical issue despite the superior performance achieved at the last layers. Unrolled networks are brittle to perturbations added to the layers’ outputs due to the lack of convergence guarantees. In contrast, standard iterative algorithms with descending guarantees, like GD, are known to be resilient to perturbations as they continue to follow descending directions in the iterations that follow the one where the perturbation occurs (see Figure 2 (left)). This resilience evidently does not extend to their unrolled counterparts since they lack descending guarantees. Revisiting our example, Figure 2 (middle) confirms this fact as the unrolled optimizer fails to converge to the same point as in Figure 1 (middle) after adding perturbations to one of its layers; in fact, the figure shows that the distance to the optimal at the last layer is larger than it is at initialization. While additive perturbations at the layers are artificial and unlikely to occur in practice, this observation indicates that unrolled networks fail to emulate the behavior of the descent algorithms they are designed to mimic. This limits our ability to draw conclusions about the generalizability of these models to out-of-distribution (OOD) problems and raises concerns about implementing them in safety-critical applications. Contrarily, standard iterative algorithms are endowed with theoretical convergence guarantees, thereby eliminating any doubts about their generalizability.
@@
Previous works have investigated the convergence issue of algorithm unrolling. One approach for providing convergence guarantees, known as safeguarding, has been proposed in [27, 28, 29, 30]. The main idea of safeguarding is to check whether the estimate made by a certain layer, during inference, satisfies some form of descent criteria. If the criteria are not satisfied, the estimate is discarded and replaced with an estimate of the classic iterative algorithm. Another approach is sequential training, a learning procedure in which the unrolled layers are trained one by one using a tailored loss function for each layer, while the previous layers either continue their training [31] or remain frozen [32]. A different line of work proved the existence of a convergent unrolled network for specific iterative algorithms such as the iterative shrinkage and thresholding algorithm (ISTA) [33, 34] and differentiable linearized ADMM [35]. However, a mechanism that guarantees to find these convergent networks was never provided. To resolve this issue, [34, 36] suggest reducing the size of the search space by learning fewer parameters of the standard algorithm, which limits the network’s expressivity. Furthermore, all these theoretical endeavors are tailored for specific optimization algorithms and do not generalize to other algorithms. This leaves many existing and potential unrolled networks without convergence guarantees.
@@
In this paper, we aim to design unrolled algorithms that follow stochastic descending directions to the optimum as in Figure 1 (right), maintain resilience to additive perturbations as illustrated Figure 2 (right) and demonstrate robustness to OOD shifts in the optimization problems they solve. On top of that, our goal is to establish theoretical convergence guarantees for unrolled networks regardless of the underlying iterative algorithms. We accomplish this by posing the training problem as a constrained learning problem. Within this framework, we enforce each unrolled layer to takes a descent direction, on average. This descent direction needs not be the gradient direction, as shown in Figure 1 (right), in order to accelerate convergence even for harder landscapes. To complement our endeavor, we theoretically derive convergence guarantees for unrolled networks trained with our descending constraints and prove exponential rates of convergence. We also show that unrolled networks maintain a descending behavior under distribution shifts, with a generalization error that depends on the distance between the original and OOD distributions. In summary, our contributions are as follows:
@@
  • We analyze the generalizability to OOD shifts and show that learning to follow a descending trajectory is transferable under OOD shifts (SectionIII-D)
  • We empirically show the performance and robustness of these unrolled networks in light of two applications: sparse coding (Section IV) and image inpainting (Section V).
@@ II Problem Formulation
The inner problem is an instantiation of an unconstrained optimization problem, which we refer to as the optimizee, where the goal is to find a minimum of $f(\cdot;{\bf x})$ for a given ${\bf x}$. It, however, needs to be evaluated for all ${\bf x}\in{\cal X}$, which is computationally expensive. The outer problem aims to learn the optimizer $\bm{\Phi}:{\cal X}\rightarrow{\cal Y}$ by minimizing a suitably-chosen loss function $\ell:{\cal Y}\times{\cal Y}\rightarrow\mathbb{R}$. For instance, the toy example in Figure 1 considers a function class $\mathfrak{F}$ that contains functions of the form $f({\bf y};{\bf x})=\|{\bf x}-{\bf Ay}\|_{2}$, which is parameterized by ${\bf x}\in{\cal X}$. The optimizee problem is then to find the stationary point of $f$ for each $\bf x$. The optimizer $\bm{\Phi}$, in this case, is designed by unrolling GD in $L$ layers, each of which is a multilayer perceptron (MLP) with residual connections, i.e., the $l$th layer’s output is ${\bf y}_{l}={\bf y}_{l-1}-\text{MLP}({\bf y}_{l-1},{\bf x};{\bf W}_{l})$. These MLPs are trained by minimzing the mean-square error (MSE) between the final prediction of $\bm{\Phi}$ and the optimal ${\bf y}^{*}$.
Solving the above bi-level optimization problem is not usually straightforward. However, it reduces to an unconstrained problem if the inner problem either has a closed-form solution or can be solved using an iterative algorithm. For our problems of interest, we assume that there exists an iterative algorithm ${\cal A}$ that finds a (local) minimum ${\bf y}^{*}$ of the optimizee problem. This algorithm is the one to be unrolled and is typically a gradient-based method, such as proximal gradient descent. Given algorithm $\cal A$ and, therefore, the optimal solution to the optimizee, the optimizer problem can be then simplified to
@@ III-A Descending Constraints
The initial value ${\bf y}_{0}$ is drawn from a Gaussian distribution ${\cal N}(\mu_{0},\sigma_{0}^{2}{\bf I})$. The expectation in the two constraints is evaluated over the joint distribution of the input $\bf x$ and the initialization ${\bf y}_{0}$. We also add noise vector ${\bf n}_{l}$, sampled from a Gaussian distribution ${\cal N}(0,\sigma^{2}_{l}{\bf I})=:{\cal N}_{l}$, to the input of layer $l$–during training–to ensure that the output ${\bf y}_{l}$ is independent of the trajectory the network follows up till layer $l$. The noise variance $\sigma_{l}^{2}$ decreases over the layers to allow the optimizer to converge.
@@
It is worth noting that the choice between (CI) or (CII) is arbitrary. The former requires the objective function $f$ to be differentiable while the latter depends on the knowledge of the optimal ${\bf y}^{*}$. Yet (CI) can be extended to non-differentiable functions by replacing the gradient with a subgradient of $f$. However, (CI) could be more challenging to satisfy if the optimizee problem is either constrained or non-convex. This is true since it could push the estimates to a non-feasible solution in the first instance or another local minimum than ${\bf y}^{*}$ in the second one. In the case of unsupervised training (as in Remark II.1), the constraints (CII) add extra computational complexity since it requires calculating the optimal solution ${\bf y}^{*}$ that is no longer required in evaluating the training loss. Other descending constraints could also be used, such as decreasing the value of the function $f$ directly.
Remark III.1.
The noise added to each layer’s output is present only during training and is omitted during execution. On one hand, it ensures that each layer takes a step independently of the path followed by the network in previous layers. On the other hand, it compels the unrolled network to explore the optimization landscape during training, which empirically helps the model’s generalizability when accompanied with the proposed descending constraints. It is pertinent to notice that our theoretical analysis in the subsequent sections remains valid when the noise is entirely omitted.
@@ III-B Probably and Approximately Correct Solution
To solve (CO), we construct the dual problem to leverage CLT [42, Theorem 1]. The dual problem of (CO) is equivalent to finding the saddle point of the Lagrangian function
@@
Solving (DO) and finding the saddle point of $\widehat{\cal L}$ is doable by alternating between minimizing with respect to ${\bf W}$ for a fixed $\bm{\lambda}$ and then maximizing the latter, as described in Algorithm 1. The convergence of this algorithm has been proved in [42, Theorem 2].
@@
These assumptions are easily achievable in practice. The Lipschitz continuity in Assumption 1 holds for a wide range of loss functions, including $\ell_{1}$ and $\ell_{2}$ norms. Modern neural networks satisfy the universal approximation theorem (Assumption 2), which demands that the family of neural networks that we consider to be sufficiently rich to approximate the update function $\phi$ up to a factor $\nu$. Assumption 3 requires the output of the neural network to be bounded, which can be ensured using a compact input set and bounded learnable parameterization. Additionally, it requires the conditional probabilities to be nonatomic, a condition that can be met by augmenting the dataset with solutions of varying precision for each instance of the optimizee problem. Assumption 4 imposes a mild assumption on the sample complexity of the models, allowing us replace the statistical expectation with the empirical average. The strict feasibility condition in Assumption 5 can also be attained by appropriately adjusting the design parameter $\epsilon$.
@@ III-C Convergence Guarantees
Given a near-optimal solution of (CO) under constraints (CI), which satisfies (8) with a probability $1-\delta$ and generates a sequence of random variables $\{{\bf y}_{l}|l\geq 0\}$. Then, under Assumption 1, it holds that
@@
$\begin{split}\mathbb{E}\big{[}{\|{\nabla}f({\bf y}_{L})\|}\big{]}\leq(1-\delta)^{L}(&1-{\epsilon})^{L}\ \mathbb{E}{\|{\nabla}f({\bf y}_{0})\|}\\ &+\frac{1}{\epsilon}\left(\zeta(N,\delta)+\frac{\delta M}{1-\delta}\right).\end{split}$ (10)
@@
Remark III.2.
Theorem 2 provides convergence guarantees for unrolled networks that parallel those of stochastic gradient-based algorithms. The similarity lies in the fact that that both algorithms follow stochastic descending directions and keep visiting a small area around the optimum infintely often. Although excursions away from this small area are possible and can be arbitrarily large, existing studies, such as [44], show that the maximum value reached in these excursions is bounded above and adheres to an exponential probability bound. These findings can be directly expanded to unrolled networks that satisfy the descending constraints.
III-D Robustness to Out-of-Distribution Shifts
The significance of the descending constraints arises from the fact that the unrolled networks learn how to find the solution by following a descending direction, rather than learning the solution itself. The latter strategy fails against OOD shifts where the input $\bf x$ is drawn from a distribution $D_{x}^{\prime}$ different from the original distribution $D_{x}$ used during training. These distribution shifts represent a change in the optimizee problem, resulting in a change in the distribution of the optimal solutions, which unrolled network cannot directly adapt to. The descending constraints, which are satisfied in expectation under the distribution $D_{x}$, can provide the unrolled networks with robustness against such distribution shifts, provided they remain satisfied under the new distribution $D_{x}^{\prime}$. The following corollaries, which hold under Assumption 6, prove that the descending nature of the layers can be transferred under distribution shifts and offer an upper bound for the generalization error. While these corollaries specifically handle the first constraints (CI),the same results can be readily expanded to the second set of constraints (CII).
Assumption 6.
There exists a non-negative symmetric distance $d$ between the input distribution $D_{x}$ and the OOD distribution $D_{x}^{\prime}$ such that
$\Big{|}\mathbb{E}_{D_{x}}\big{[}{\|{\nabla}f({\bf y};{\bf x})\|}\Big{]}-\mathbb{E}_{D_{x}^{\prime}}\big{[}{\|{\nabla}f({\bf y};{\bf x})\|}\big{]}\Big{|}\leq Md(D_{x},D_{x}^{\prime})$
uniformly over ${\bf y}$ with $M$ being a Lipschitz constant.
This assumption is easily attainable. For instance, the Wasserstein distance $W_{1}(D_{x},D_{x}^{\prime})$ satisfies this assumption and provides a bounded distance between the two distributions $D_{x}$ and $D_{x}^{\prime}$. Given an unrolled network trained on the distribution $D_{x}$, the following corollary evaluates the descending constraints when the network is executed under the distribution $D_{x}^{\prime}$.
Corollary 1.
Under Assumption 6, an unrolled network ${\bf W}$ trained via Algorithm 1 over a distribution $D_{x}$ generates when executed on an OOD distribution $D_{x}^{\prime}$ a sequence of layers’ outputs $\{{\bf y}_{l}\}_{l}$ that satisfies
$\begin{split}\mathbb{E}_{D_{x}^{\prime}}\big{[}\|{\nabla}f({\bf y}_{l};{\bf x})\|_{2}\ -&\ (1-\epsilon)\ \|{\nabla}f({\bf y}_{l-1};{\bf x})\|_{2}\big{]}\\ &\leq\zeta(N,\delta)+2Md(D_{x},D_{x}^{\prime}),\end{split}$ (11)
with probability $1-\delta$ each.
Proof.
The proof can be found in Appendix C. ∎
Corollary 1 indicates that the descending constraints are satisfied under an OOD distribution $D_{x}^{\prime}$ up to a generalization error that depends on the distance between the two distributions $D_{x}$ and $D_{x}^{\prime}$. Even under distribution shifts, the unrolled network still finds (noisy) descending directions toward the optimum, despite not being trained on the solutions of the new optimizee problems. This capability is attributed to the proposed descending constraints and is unique to the unrolled networks trained with such constraints.
Since the descending constraints are satisfied with a marginal error and with probability $1-\delta$, convergence guarantees for the unrolled network can be derived under the new distribution similarly to Theorem 2. The following corollary summarizes this result.
Corollary 2.
The sequence of layers’ outputs $\{{\bf y}_{l}\}_{l}$ generated according to Corollary 1 also satisfies
$\begin{split}\lim_{l\rightarrow\infty}\mathbb{E}_{D_{x}^{\prime}}\Big{[}&\min_{k\leq l}\|{\nabla}f({\bf y}_{k};{\bf x})\|\Big{]}\\ &\leq\frac{1}{\epsilon}\Big{(}\zeta(N,\delta)+2Md(D_{x},D_{x}^{\prime})+\frac{\delta M}{1-\delta}\Big{)}.\end{split}$ (12)
This corollary follows directly from Theorem 2. It states that, under distribution shifts, the unrolled network converges eventually to a region around the optimal solution where the gradient norm is below a factor that includes the distance between the two distributions. The smaller this distance, the smaller the near-optimal region to which the network converges.
@@ IV-A Problem Formulation
$\begin{split}\min_{{\bf y}}\quad&f_{sp}({\bf y};{\bf x}):=\frac{1}{2}\|{\bf x}-{\bf Dy}\|_{2}^{2}+\alpha\|{\bf y}\|_{1},\end{split}$ (13)
@@
Figure 3: Distance to the optimal solution ${\bf y}^{*}$ and the value of the objective function $f_{sp}({\bf y}_{l};{\bf x})$ across the ten unrolled layers of constrained LISTA (blue), LISTA (red), incremental training (black), and layer-per-layer training (black dots). Constrained LISTA makes gradual progressions toward the optimum, unlike the other approaches, due to the implemented descending constraints during training.
@@ IV-D Numerical Results and Discussions
Performance. Figure 3 compares the performance of LISTA and constrained LISTA on a test dataset. Both models achieve equivalent performance at their outputs, illustrated by the zero distance achieved at the last layer in both cases. However, their behaviors over the intermediate layers differ significantly from each other. For standard LISTA, the distance to the optimal is almost the same across all these layers, which suggests that there was no significant improvement in their estimates. This is further confirmed by examining the value of the objective function $f_{sp}({\bf y},{\bf x})$ in Figure 3 (right), which shows negligible change across the intermediate layers. A similar pattern is observed for other training procedures, such as incremental training [31] and layer-per-layer training [32]. In contrast, constrained LISTA demonstrates a gradual decrease in the distance to the optimum and the objective function because of the imposed constraints.
Figure 5: OOD robustness against data shifts of the form $\tilde{\bf x}\sim{\cal N}({\bf x},p^{2}{\bf I})$. Distance to the optimal solution $\tilde{\bf y}^{*}$ and the value of the objective function $f_{sp}({\bf y}_{l})$ across the ten unrolled layers of constrained LISTA (blue), LISTA (red), LISTA trained with noisy inputs (gray), and other benchmarks (black).
@@
These observations suggest that standard LISTA can find the solution in just two layers, where a significant reduction in the distance to the optimum is observed. In contrast, constrained LISTA requires more layers for a gradual reduction to occur. In order to examine whether we can combine the advantages of the two models–namely, achieving substantial progress in the first layer and then allowing the network to descend–we train LISTA with constraints starting from the second layer, leaving the first layer unconstrained. Figure 3 shows that the model prioritizes learning a descending path over the constrained layers. However, unlike standard unrolling, the unconstrained first layer follows an ascending direction, where both the distance to the optimum and the the value of $f_{sp}$ increase. This implies that the direction predicted by an unconstrained layer is largely arbitrary.
Robustness. To evaluate the impact of the descending constraints, we test the unrolled network on an OOD distribution. The OOD dataset $\{(\tilde{\bf x}_{i},\tilde{\bf y}_{i}^{*})\}_{i}$ is constructed by adding perturbations to the input data ${\bf x}\sim D_{x}$ and then running ISTA from scratch to find the optimal sparse representations of the new signals $\tilde{\bf x}$. The perturbations are randomly drawn from a Gaussian distribution ${\cal N}(0,p^{2}{\bf I})$ and the larger the level of perturbations $p$, the larger the distance $d(D_{x},D_{x}^{\prime})$. Figure 5 shows the robustness of constrained LISTA against the distribution shift, as indicated by the slow increase in the distance to the optimal as the perturbation size increases. In contrast, standard LISTA and the other benchmarks experience a higher distance to the optimum and a higher value of the objective function for all values of $p$. Additionally, the gap between the two models widen significantly with the increase in perturbation size.
Figure 6: Robustness against additive noise with a standard deviation proportional to the gradient by a factor $\hat{\sigma}$ (training is held at $\hat{\sigma}=1$). (Left) Distance to the optimal ${\bf y}^{*}$, (middle) $\ell_{1}$-norm of the unrolled network’s output, and (right) Frequency of satisfying the constraints across the test dataset for different perturbation sizes.
To ensure a fair comparison, we also train standard LISTA with the same additive noise schedule introduced in Section III-A. Although adding perturbations during training might suggest enhanced robustness of the model, Figure 5 shows that incorporating noise during training does not significantly improve the generalizability of standard LISTA. Furthermore, Figure 6 illustrates the performance of both LISTA and constrained LISTA when subjected to additive perturbations at each unrolled layer during execution, similar to the noise introduced during training but with different magnitudes. The figure confirms that training the models with noise does confer some robustness toward this type of perturbations. However, noisy training alone in insufficient to ensure robustness toward distribution shifts in the input data. This finding confirms that the OOD robustness observed in constrained LISTA, as shown in Figure 5, is entirely attributed to the proposed descending constraints and the convergence guarantees they provide, even under distribution shifts.